Circular No. 6025

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Further reports received via the SL9 message center include detections

of the impacts on Jupiter of fragments D = 18 and E = 17.

CASPIR 2.34-μm images and IRIS K-grism cubes at Siding Spring (cf. IAUC 6024) revealed a faint flash due to the D impact on July 17.496 UT that faded within seconds. A second flash on July 17.500 was about four times brighter than the C impact site (then on the opposite limb) and about half as bright as Europa in the K-band, but it faded in < 1 min. The K-grism results then revealed a slow brightening during July 17.512–17.521 and a subsequent decay. CASPIR 3.28- and 3.42-μm images taken near July 17.514 indicated that the feature was somewhat brighter at these wavelengths. There was no evidence of the impact in IRIS H-grism or CASPIR 4.78-μm images.

J. Watanabe and others report the appearance of the plume of fragment D on July 17.501 UT in 2.36-μm images obtained with the 1.88-m telescope at the Okayama Astrophysical Observatory. It was fainter than the plume recorded from the impact of C on July 17.303; that plume was brightest around July 17.306, and the decay phase was observed over one hour.

The French-Swedish-Spanish Nordic Optical Telescope team (cf. IAUC 6023) reports that the impact of fragment E was recorded at 10 μm around July 17.637 UT. It was fainter than impact A seen yesterday.

The Calar Alto Observing Team (cf. IAUC 6023) reports the detection of the fireball from fragment E with the 3.5-m Calar Alto telescope in the 2.3-μm band. The event was first noted on July 17.637 UT, and it rapidly increased to > 30 times the brightness of Europa. By July 17.641 the fireball had dimmed to less than Europa’s brightness. On July 17.768 three spots were visible—a bright one (impact E) just past the central meridian, a fainter one (impact A) approaching the central meridian and the faintest one (impact C) just coming over the east limb. At 2.3 μm the brightest spot was 15–20 times fainter than Io, and the others were fainter by further factors of ∼ 1.5 and 2.2.

J. Rogers, British Astronomical Association, reports his visual detection (before sunset using yellow and polaroid filters) with the 0.30-m refractor at the University of Cambridge of conspicuous dark spots at the predicted impact sites of fragments A and C. Site A, which has clearly grown since the first Hubble image, was the darker (similar to the NEB) and remained visible until it approached Jupiter’s limb on July 17.851 UT. Site C was on the central meridian on July 17.828 ± 0.002.

1994 July 17 Brian G. Marsden